Acounting

A machine costing $210,000 with a four-year life and an estimated $20,000 salvage value is installed in Calhoon Company’s factory on January 1. The factory managerestimates the machine will produce 475,000 units of product during its life. It actually produces the following units: year 1, 121,400; year 2, 122,400; year 3,119,600; and year 4, 118,200. The total number of units produced by the end of year 4 exceeds the original estimate—this difference was not predicted. (The machinemust not be depreciated below its estimated salvage value.)
Answer #1

Straight-Line (210,000- 20,000)/4 =47,500 Units of Production 210,000 - 20,000 = $190,000 Depreciable amount (210,000- 20,000) / 475,000 = $0.40 perunit Year One is 121,400 x 0.40 = 48,560 So year 2 should be 122,400 x 0.40 = 48,960 Since you're not allowed to depreciate more than the depreciableamount, year 4 would be 190,000 - 48,560 - 48,960 - 47,840 = $44,640 Double-Declining Balance 100% / 4 year useful life = 25% straight-linedepreciation rate 25% x 2 = 50% double-declining rate Year 1 210,000 x 50% = $105,000 depreciation expense Year 2 105,000 x 50% = $52,000 depreciationexpense Year 3 52,000 x 50% = $26,250 depreciation expense Year 4 *Note: When using the double-declining method, the book value is not allowed to dropbelow the salvage value of the asset. The depreciation expense would be 26,250-22,000 = $6,250.

answered by: Zaki
Answer #2
Part A
Year : Straight-Line; Units-of-Production; Double-Declining-Balance
1 : $47,500; $48,560; $105,000
2 : $47,500; $48,960; $52,500
3 : $47,500; $47,840; $26,250
4 : $47,500; $44,640; $6,250


answered by: Maxy